# Finite Sample Analysis of Distributional TD Learning with Linear Function Approximation

### v1

@@ C.2 Linear-Categorical Projection Operator
$=2\iota_{K}\left({{\bm{I}}_{K}\otimes{\bm{\phi}}(s)}\right){\bm{C}}^{\intercal}{\bm{C}}\left({{\bm{p}}_{{\bm{w}}}(s)-{\bm{p}}_{\bm{\eta}}(s)}\right)$
$=2\iota_{K}\left({{\bm{I}}_{K}\otimes{\bm{\phi}}(s)}\right){\bm{C}}^{\intercal}{\bm{C}}\left({\left({{\bm{I}}_{K}\otimes{\bm{\phi}}(s)}\right)^{\intercal}{\bm{w}}+\frac{1}{K+1}\bm{1}_{K}-{\bm{p}}_{\bm{\eta}}(s)}\right)$
$=2\iota_{K}\left\{{\left[{\left({{\bm{C}}^{\intercal}{\bm{C}}}\right)\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\intercal}}\right)}\right]{\bm{w}}+\left[{\left({{\bm{C}}^{\intercal}{\bm{C}}\left({\frac{1}{K+1}\bm{1}_{K}-{\bm{p}}_{\bm{\eta}}(s)}\right)}\right)\otimes{\bm{\phi}}(s)}\right]}\right\},$
where in the second equality, we used Proposition B.1, and in the last equality, we used Lemma A.5. We also have the following matrix representation:
$\nabla_{{\bm{W}}}\ell_{2}^{2}\left({{\eta}_{{\bm{w}}}(s),{\eta}(s)}\right)$ $=2\iota_{K}{\bm{\phi}}(s)\left({{\bm{p}}_{{\bm{w}}}(s)-{\bm{p}}_{\bm{\eta}}(s)}\right)^{\intercal}{\bm{C}}^{\intercal}{\bm{C}}$
$=2\iota_{K}{\bm{\phi}}(s)\left[{{\bm{\phi}}(s)^{\intercal}{\bm{W}}+\left({\frac{1}{K+1}\bm{1}_{K}-{\bm{p}}_{\bm{\eta}}(s)}\right)^{\intercal}}\right]{\bm{C}}^{\intercal}{\bm{C}}.$
Proposition C.2.
For any $\bm{\eta}\in\left({{\mathscr{P}}^{\mathsf{sign}}}\right)^{{\mathcal{S}}}$ and $\bm{\eta}_{{\bm{w}}}\in{\mathscr{P}}^{\mathsf{sign}}_{{\bm{\phi}},K}$, it holds that
$\ell_{2,\mu_{\pi}}^{2}\left({\bm{\eta},\bm{\eta}_{{\bm{w}}}}\right)=\ell_{2,\mu_{\pi}}^{2}\left({\bm{\eta},{\bm{\Pi}}_{K}\bm{\eta}}\right)+\ell_{2,\mu_{\pi}}^{2}\left({{\bm{\Pi}}_{K}\bm{\eta},{\bm{\Pi}}_{{\bm{\phi}},K}^{\pi}\bm{\eta}}\right)+\ell_{2,\mu_{\pi}}^{2}\left({{\bm{\Pi}}_{{\bm{\phi}},K}^{\pi}\bm{\eta},\bm{\eta}_{{\bm{w}}}}\right).$
The proof is straightforward and almost the same as that of Proposition B.2 if we utilize the affine structure.
C.3 Linear-Categorical Projected Bellman Equation
To derive the result, the following proposition characterizing ${\bm{\Pi}}_{{\bm{\phi}},K}^{\pi}\bm{\mathcal{T}}^{\pi}\bm{\eta}_{{\bm{w}}}$ is useful, whose proof can be found in Appendix C.4.
Proposition C.3.
For any ${\bm{w}}\in{\mathbb{R}}^{dK}$ and $s\in{\mathcal{S}}$, we abbreviate ${\bm{p}}_{\bm{\mathcal{T}}^{\pi}\bm{\eta}_{\bm{w}}}(s)$ as $\tilde{{\bm{p}}}_{{\bm{w}}}(s)$, then
$\tilde{{\bm{p}}}_{{\bm{w}}}(s)=$ $\left({\tilde{p}_{k}(s;{\bm{w}})}\right)_{k=0}^{K-1}={\mathbb{E}}\left[{\tilde{{\bm{G}}}(r_{0}){\bm{W}}^{\intercal}{\bm{\phi}}(s_{1})\Big{|}s_{0}=s}\right]+\frac{1}{K+1}\sum_{j=0}^{K}{\mathbb{E}}\left[{{\bm{g}}_{j}(r_{0})\Big{|}s_{0}=s}\right],$
Combining this proposition with Proposition 3.1, we know that ${\bm{W}}^{\star}$ is the unique solution to the following system of linear equations for ${\bm{W}}\in{\mathbb{R}}^{d\times K}$
${\bm{W}}=$ ${\bm{\Sigma}}_{{\bm{\phi}}}^{-1}{\mathbb{E}}_{s\sim\mu_{\pi}}\left[{{\bm{\phi}}(s)\left({\tilde{\bm{p}}_{{\bm{w}}}(s)-\frac{1}{K+1}\bm{1}_{K}}\right)^{\intercal}}\right]$
$=$ ${\bm{\Sigma}}_{{\bm{\phi}}}^{-1}{\mathbb{E}}_{s\sim\mu_{\pi}}\left[{{\bm{\phi}}(s)\left({\frac{1}{K+1}\bm{1}_{K}-{\mathbb{E}}\left[{\tilde{{\bm{G}}}(r_{0}){\bm{W}}^{\intercal}{\bm{\phi}}(s_{1})\Big{|}s_{0}=s}\right]-\frac{1}{K+1}\sum_{j=0}^{K}{\mathbb{E}}\left[{{\bm{g}}_{j}(r_{0})\Big{|}s_{0}=s}\right]}\right)^{\intercal}}\right]$
$=$ ${\bm{\Sigma}}_{{\bm{\phi}}}^{-1}{\mathbb{E}}_{s\sim\mu_{\pi}}\left[{{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}{\bm{W}}\tilde{{\bm{G}}}^{\intercal}(r)}\right]+\frac{1}{K+1}{\bm{\Sigma}}_{{\bm{\phi}}}^{-1}{\mathbb{E}}_{s\sim\mu_{\pi}}\left[{{\bm{\phi}}(s)\left({\sum_{j=0}^{K}{\bm{g}}_{j}(r)-\bm{1}_{K}}\right)^{\intercal}}\right],$
or equivalently,
${\bm{\Sigma}}_{{\bm{\phi}}}{\bm{W}}-{\mathbb{E}}_{s\sim\mu_{\pi}}\left[{{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}{\bm{W}}\tilde{{\bm{G}}}^{\intercal}(r)}\right]=\frac{1}{K+1}{\mathbb{E}}_{s\sim\mu_{\pi}}\left[{{\bm{\phi}}(s)\left({\sum_{j=0}^{K}{\bm{g}}_{j}(r)-\bm{1}_{K}}\right)^{\intercal}}\right],$
which is the desired conclusion. The uniqueness and existence of the solution is guaranteed by the fact that the LHS is an invertible linear transformation of ${\bm{W}}$, which is justified by Eqn. (25).
C.4 Proof of Proposition C.3
Proof.
$\tilde{p}_{k}(s;{\bm{w}})=$ $p_{k}\left({\left[{\bm{\mathcal{T}}^{\pi}\bm{\eta}_{\bm{w}}}\right](s)}\right)$ (17)
$=$ ${\mathbb{E}}_{X\sim\left[{\bm{\mathcal{T}}^{\pi}\bm{\eta}_{\bm{w}}}\right](s)}\left[{\left({1-\left|{\frac{X-x_{k}}{\iota_{K}}}\right|}\right)_{+}}\right]$
$=$ ${\mathbb{E}}\left[{{\mathbb{E}}_{G\sim\eta_{{\bm{w}}}(s_{1})}\left[{\left({1-\left|{\frac{r_{0}+\gamma G-x_{k}}{\iota_{K}}}\right|}\right)_{+}}\right]\Big{|}s_{0}=s}\right]$
$=$ ${\mathbb{E}}\left[{\sum_{j=0}^{K}p_{j}(s_{1};{\bm{w}})\left({1-\left|{\frac{r_{0}+\gamma x_{j}-x_{k}}{\iota_{K}}}\right|}\right)_{+}\Big{|}s_{0}=s}\right]$
$=$ ${\mathbb{E}}\left[{\sum_{j=0}^{K}p_{j}(s_{1};{\bm{w}})g_{j,k}(r_{0})\Big{|}s_{0}=s}\right]$
$=$ ${\mathbb{E}}\left[{g_{K,k}(r_{0})+\sum_{j=0}^{K-1}p_{j}(s_{1};{\bm{w}})\left({g_{j,k}(r_{0})-g_{K,k}(r_{0})}\right)\Big{|}s_{0}=s}\right].$
Hence,
$\tilde{{\bm{p}}}_{{\bm{w}}}(s)=$ $\left({\tilde{p}_{k}(s;{\bm{w}})}\right)_{k=0}^{K-1}$
$=$ ${\mathbb{E}}\left[{\begin{bmatrix}g_{K,1}(r_{0})\\ \vdots\\ g_{K,{K-1}}(r_{0})\end{bmatrix}+\sum_{j=0}^{K-1}p_{j}(s_{1};{\bm{w}})\begin{bmatrix}g_{j,1}(r_{0})-g_{K,1}(r_{0})\\ \vdots\\ g_{j,K-1}(r_{0})-g_{K,K-1}(r_{0})\end{bmatrix}\Bigg{|}s_{0}=s}\right]$
$=$ ${\mathbb{E}}\left[{{\bm{g}}_{K}(r_{0})+\sum_{j=0}^{K-1}p_{j}(s_{1};{\bm{w}})\left({{\bm{g}}_{j}(r_{0})-{\bm{g}}_{K}(r_{0})}\right)\Big{|}s_{0}=s}\right]$
$=$ ${\mathbb{E}}\left[{{\bm{g}}_{K}(r_{0})+\left({{\bm{G}}(r_{0})-\bm{1}_{K}^{\intercal}\otimes{\bm{g}}_{K}(r_{0})}\right){\bm{p}}_{{\bm{w}}}(s_{1})\Big{|}s_{0}=s}\right]$
$=$ ${\mathbb{E}}\left[{{\bm{g}}_{K}(r_{0})+\left({{\bm{G}}(r_{0})-\bm{1}_{K}^{\intercal}\otimes{\bm{g}}_{K}(r_{0})}\right)\left[{\left({{\bm{I}}_{K}\otimes{\bm{\phi}}(s_{1})}\right)^{\intercal}{\bm{w}}+\frac{1}{K+1}\bm{1}_{K}}\right]\Big{|}s_{0}=s}\right]$
$=$ ${\mathbb{E}}\left[{\left({{\bm{G}}(r_{0})-\bm{1}_{K}^{\intercal}\otimes{\bm{g}}_{K}(r_{0})}\right)\left({{\bm{I}}_{K}\otimes{\bm{\phi}}(s_{1})}\right)^{\intercal}\Big{|}s_{0}=s}\right]{\bm{w}}$
$\quad+{\mathbb{E}}\left[{{\bm{g}}_{K}(r_{0})+\frac{1}{K+1}\left({{\bm{G}}(r_{0})-\bm{1}_{K}^{\intercal}\otimes{\bm{g}}_{K}(r_{0})}\right)\bm{1}_{K}\Big{|}s_{0}=s}\right]$
$=$ ${\mathbb{E}}\left[{\left({{\bm{G}}(r_{0})-\bm{1}_{K}^{\intercal}\otimes{\bm{g}}_{K}(r_{0})}\right)\otimes{\bm{\phi}}(s_{1})^{\intercal}\Big{|}s_{0}=s}\right]{\bm{w}}+\frac{1}{K+1}\sum_{j=0}^{K}{\mathbb{E}}\left[{{\bm{g}}_{j}(r_{0})\Big{|}s_{0}=s}\right],$
or equivalently,
$\tilde{{\bm{p}}}_{{\bm{w}}}(s)={\mathbb{E}}\left[{\tilde{{\bm{G}}}(r_{0}){\bm{W}}^{\intercal}{\bm{\phi}}(s_{1})\Big{|}s_{0}=s}\right]+\frac{1}{K+1}\sum_{j=0}^{K}{\mathbb{E}}\left[{{\bm{g}}_{j}(r_{0})\Big{|}s_{0}=s}\right].$
C.5 Proof of Proposition 3.2
Proof.
By the basic inequality (Lemma E.1), we only need to show
$\ell_{2,\mu_{\pi}}^{2}\left({\bm{\eta}^{\pi},\bm{\eta}_{{\bm{w}}^{\star}}}\right)\leq$ $\frac{\ell_{2,\mu_{\pi}}^{2}\left({\bm{\eta}^{\pi},{\bm{\Pi}}_{{\bm{\phi}},K}^{\pi}\bm{\eta}^{\pi}}\right)}{1-\gamma}$
where we used [Proposition 9.18 and Eqn. (5.28) Bellemare et al., 2023]. ∎
C.6 Cumulative Distribution Function Representation
C.6.1 Categorical Parametrization
For any $\nu\in{\mathscr{P}}^{\mathsf{sign}}_{K}$, we denote
${\bm{F}}_{\nu}=\left({F_{k}\left({\nu}\right)}\right)_{k=0}^{K-1}=\left({\nu\left({[0,x_{k}]}\right)}\right)_{k=0}^{K-1}\in{\mathbb{R}}^{K}$
as the cumulative distribution function (CDF) representation of $\nu$. One can check that ${\bm{F}}_{\nu}={\bm{C}}{\bm{p}}_{\nu}$. As a result, PMF and CDF representations are equivalent because ${\bm{C}}$ is invertible.
For any $\nu_{1},\nu_{2}\in{\mathscr{P}}^{\mathsf{sign}}_{K}$,
$\ell_{2}^{2}(\nu_{1},\nu_{2})=$ $\iota_{K}\left\|{{\bm{F}}_{\nu_{1}}-{\bm{F}}_{\nu_{2}}}\right\|^{2},$
therefore, the affine space $\left({{\mathscr{P}}^{\mathsf{sign}}_{K},\ell_{2}}\right)$ is isometric with the Euclidean space $\left({{\mathbb{R}}^{K},\sqrt{\iota_{K}}\left\|{\cdot}\right\|_{2}}\right)$ if we consider the CDF representation.
For any ${\bm{\eta}}\in\left({{\mathscr{P}}^{\mathsf{sign}}_{K}}\right)^{\mathcal{S}}$, $s\in{\mathcal{S}}$, we abuse the notation to define ${\bm{F}}_{\bm{\eta}}(s):={\bm{F}}_{\eta(s)}$. By Eqn. (5), we can also similarly define ${\bm{F}}_{\nu},{\bm{F}}_{\bm{\eta}}(s)$ using the categorical projection operator for any $\nu\in{\mathscr{P}}^{\mathsf{sign}}$, $\bm{\eta}\in\left({{\mathscr{P}}^{\mathsf{sign}}}\right)^{{\mathcal{S}}}$ and $s\in{\mathcal{S}}$.
C.6.2 Linear-Categorical Parametrization
We introduce new notations for the CDF representation when we consider linear-categorical parametrization. Let ${\bm{\Theta}}:={\bm{W}}{\bm{C}}^{\intercal}=\left({\sum_{k=0}^{0}{\bm{w}}(k),\sum_{k=0}^{1}{\bm{w}}(k),\cdots,\sum_{k=0}^{K}{\bm{w}}(k)}\right)\in{\mathbb{R}}^{d\times K}$ and the vectorization of ${\bm{\Theta}}$, ${\bm{\theta}}:=\mathsf{vec}\left({{\bm{\Theta}}}\right)=\left({{\bm{C}}\otimes{\bm{I}}_{d}}\right){\bm{w}}\in{\mathbb{R}}^{dK}$. We abbreviate ${\bm{F}}_{\bm{\eta}_{\bm{w}}}$ as ${\bm{F}}_{{\bm{\theta}}}$, then by Lemma A.2 and Lemma A.5, for any $s\in{\mathcal{S}}$, it holds that
${\bm{F}}_{{\bm{\theta}}}(s)={\bm{C}}{\bm{p}}_{{\bm{w}}}(s)={\bm{\Theta}}^{\intercal}{\bm{\phi}}(s)+\frac{1}{K+1}{\bm{C}}\bm{1}_{K}=\left({{\bm{I}}_{K}\otimes{\bm{\phi}}(s)}\right)^{\intercal}{\bm{\theta}}+\frac{1}{K+1}{\bm{C}}\bm{1}_{K},$ (18)
Again, PMF and CDF representations are equivalent because ${\bm{C}}$ is invertible.
For any $\bm{\eta}_{{\bm{w}}_{1}},\bm{\eta}_{{\bm{w}}_{2}}\in{\mathscr{P}}^{\mathsf{sign}}_{{\bm{\phi}},K}$, by Proposition B.1,
$=$ $\iota_{K}{\mathbb{E}}_{s\sim\mu_{\pi}}\left[{\left\|{{\bm{F}}_{{\bm{\theta}}_{1}}(s)-{\bm{F}}_{{\bm{\theta}}_{2}}(s)}\right\|^{2}}\right]$
$=$ $\iota_{K}\operatorname{tr}\left({\left({{\bm{\Theta}}_{1}-{\bm{\Theta}}_{2}}\right)^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}\left({{\bm{\Theta}}_{1}-{\bm{\Theta}}_{2}}\right)}\right)$
$=$ $\iota_{K}\left\|{{\bm{\theta}}_{1}-{\bm{\theta}}_{2}}\right\|^{2}_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}},$
hence the affine space $\left({{\mathscr{P}}^{\mathsf{sign}}_{{\bm{\phi}},K},\ell_{2,\mu_{\pi}}}\right)$ is isometric with the Euclidean space $\left({{\mathbb{R}}^{Kd},\sqrt{\iota_{K}}\left\|{\cdot}\right\|_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}}\right)$ if we consider the CDF representation.
Following the proof of Lemma C.1 in Appendix C.2, we can also derive the gradient when we use the CDF parametrization:
$\nabla_{{\bm{\theta}}}\ell_{2}^{2}\left({{\eta}_{{\bm{w}}}(s),{\eta}(s)}\right)$ $=\iota_{K}\nabla_{{\bm{\theta}}}\left\|{{\bm{F}}_{{\bm{\theta}}}(s)-{\bm{F}}_{\bm{\eta}}(s)}\right\|^{2}$
$=2\iota_{K}\left({{\bm{I}}_{K}\otimes{\bm{\phi}}(s)}\right)\left({\left({{\bm{I}}_{K}\otimes{\bm{\phi}}(s)}\right)^{\intercal}{\bm{\theta}}+{\bm{C}}\left({\frac{1}{K+1}\bm{1}_{K}-{\bm{p}}_{\bm{\eta}}(s)}\right)}\right)$
$=2\iota_{K}\left\{{\left[{{\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\intercal}}\right)}\right]{\bm{\theta}}+\left[{\left({{\bm{C}}\left({\frac{1}{K+1}\bm{1}_{K}-{\bm{p}}_{\bm{\eta}}(s)}\right)}\right)\otimes{\bm{\phi}}(s)}\right]}\right\},$
$\nabla_{{\bm{\Theta}}}\ell_{2}^{2}\left({{\eta}_{{\bm{w}}}(s),{\eta}(s)}\right)$ $=2\iota_{K}{\bm{\phi}}(s)\left({{\bm{F}}_{{\bm{\theta}}}(s)-{\bm{F}}_{\bm{\eta}}(s)}\right)^{\intercal}$ (20)
$=2\iota_{K}{\bm{\phi}}(s)\left[{{\bm{\phi}}(s)^{\intercal}{\bm{\Theta}}+\left({\frac{1}{K+1}\bm{1}_{K}-{\bm{p}}_{\bm{\eta}}(s)}\right)^{\intercal}{\bm{C}}^{\intercal}}\right].$
C.7 Stochastic Semi-Gradient Descent with Linear Function Approximation
We use the notations in the generative model setting for simplicity. We denote by ${\bm{\mathcal{T}}}_{t}^{\pi}$ the corresponding empirical distributional Bellman operator at the $t$-th iteration, which satisfies
$\left[{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}}}\right](s_{t})=(b_{r_{t},\gamma})_{\#}(\eta(s_{t}^{\prime})),\quad\forall\bm{\eta}\in{\mathscr{P}}^{{\mathcal{S}}}.$ (21)
C.7.1 PMF Representation
Consider the SSGD with the PMF representation
${\bm{W}}_{t}={\bm{W}}_{t-1}-\alpha\nabla_{{\bm{W}}}\ell_{2}^{2}\left({{\eta}_{{\bm{w}}_{t-1}}(s_{t}),\left[{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}_{{\bm{w}}_{t-1}}}}\right](s_{t})}\right),$
where $\nabla_{{\bm{W}}}$ stands for taking gradient w.r.t. ${\bm{W}}_{t-1}\in{\mathbb{R}}^{d\times K}$ in the first term ${\eta}_{{\bm{w}}_{t-1}}(s_{t})$ (the second term is regarding as a constant, that’s why we call it a semi-gradient). One can check that $\nabla_{{\bm{W}}}\ell_{2}^{2}\left({{\eta}_{{\bm{w}}_{t-1}}(s_{t}),\left[{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}_{{\bm{w}}_{t-1}}}}\right](s_{t})}\right)$ is an unbiased estimate of $\nabla_{{\bm{W}}}\ell_{2,\mu_{\pi}}^{2}\left({\bm{\eta}_{{\bm{w}}_{t-1}},\bm{\mathcal{T}}^{\pi}{\bm{\eta}_{{\bm{w}}_{t-1}}}}\right)$.
Now, let’s compute the gradient term. By Lemma C.1, we have
$\nabla_{{\bm{W}}}\ell_{2}^{2}\left({{\eta}_{{\bm{w}}_{t-1}}(s_{t}),\left[{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}_{{\bm{w}}_{t-1}}}}\right](s_{t})}\right)$ $=2\iota_{K}{\bm{\phi}}(s_{t})\left({{\bm{p}}_{{\bm{w}}_{t-1}}(s_{t})-{\bm{p}}_{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}_{{\bm{w}}_{t-1}}}}(s_{t})}\right)^{\intercal}{\bm{C}}^{\intercal}{\bm{C}}$
$=2\iota_{K}{\bm{\phi}}(s_{t})\left[{{\bm{\phi}}(s_{t})^{\intercal}{\bm{W}}_{t-1}+\left({\frac{1}{K+1}\bm{1}_{K}-{\bm{p}}_{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}_{{\bm{w}}_{t-1}}}}(s_{t})}\right)^{\intercal}}\right]{\bm{C}}^{\intercal}{\bm{C}}.$

### v2

@@ D.6 Linear-Categorical Parametrization with Probability Mass Function Representation
$=2\iota_{K}\left({{\bm{I}}_{K}\otimes{\bm{\phi}}(s)}\right){\bm{C}}^{\top}{\bm{C}}\left({{\bm{p}}_{{\bm{w}}}(s)-{\bm{p}}_{\bm{\eta}}(s)}\right)$
$=2\iota_{K}\left({{\bm{I}}_{K}\otimes{\bm{\phi}}(s)}\right){\bm{C}}^{\top}{\bm{C}}\left({\left({{\bm{I}}_{K}\otimes{\bm{\phi}}(s)}\right)^{\top}{\bm{w}}+\frac{1}{K+1}\bm{1}_{K}-{\bm{p}}_{\bm{\eta}}(s)}\right)$
$=2\iota_{K}\left\{{\left[{\left({{\bm{C}}^{\top}{\bm{C}}}\right)\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\top}}\right)}\right]{\bm{w}}+\left[{\left({{\bm{C}}^{\top}{\bm{C}}\left({\frac{1}{K+1}\bm{1}_{K}-{\bm{p}}_{\bm{\eta}}(s)}\right)}\right)\otimes{\bm{\phi}}(s)}\right]}\right\},$
$\nabla_{{\bm{W}}}\ell_{2}^{2}\left({{\eta}_{{\bm{w}}}(s),{\eta}(s)}\right)$ $=2\iota_{K}{\bm{\phi}}(s)\left({{\bm{p}}_{{\bm{w}}}(s)-{\bm{p}}_{\bm{\eta}}(s)}\right)^{\top}{\bm{C}}^{\top}{\bm{C}}$ (23)
$=2\iota_{K}{\bm{\phi}}(s)\left[{{\bm{\phi}}(s)^{\top}{\bm{W}}+\left({\frac{1}{K+1}\bm{1}_{K}-{\bm{p}}_{\bm{\eta}}(s)}\right)^{\top}}\right]{\bm{C}}^{\top}{\bm{C}}.$
D.7 Stochastic Semi-Gradient Descent with Linear Function Approximation
We denote by ${\bm{\mathcal{T}}}_{t}^{\pi}$ the corresponding empirical distributional Bellman operator at the $t$-th iteration, which satisfies
$\left[{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}}}\right](s_{t})=(b_{r_{t},\gamma})_{\#}(\eta(s_{t+1})),\quad\forall\bm{\eta}\in{\mathscr{P}}^{{\mathcal{S}}}.$ (24)
D.7.1 Probability Mass Function Representation
Consider the stochastic semi-gradient descent (SSGD) with the probability mass function (PMF) representation
${\bm{W}}_{t}\leftarrow{\bm{W}}_{t-1}-\alpha\nabla_{{\bm{W}}}\ell_{2}^{2}\left({{\eta}_{{\bm{w}}_{t-1}}(s_{t}),\left[{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}_{{\bm{w}}_{t-1}}}}\right](s_{t})}\right),$
where $\nabla_{{\bm{W}}}$ stands for taking gradient w.r.t. ${\bm{W}}_{t-1}\in{\mathbb{R}}^{d\times K}$ in the first term ${\eta}_{{\bm{w}}_{t-1}}(s_{t})$ (the second term is regarding as a constant, that’s why we call it a semi-gradient). We can check that $\nabla_{{\bm{W}}}\ell_{2}^{2}\left({{\eta}_{{\bm{w}}_{t-1}}(s_{t}),\left[{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}_{{\bm{w}}_{t-1}}}}\right](s_{t})}\right)$ is an unbiased estimate of $\nabla_{{\bm{W}}}\ell_{2,\mu_{\pi}}^{2}\left({\bm{\eta}_{{\bm{w}}_{t-1}},\bm{\mathcal{T}}^{\pi}{\bm{\eta}_{{\bm{w}}_{t-1}}}}\right)$.
Now, let us compute the gradient term. By Eqn. (23), we have
$\nabla_{{\bm{W}}}\ell_{2}^{2}\left({{\eta}_{{\bm{w}}_{t-1}}(s_{t}),\left[{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}_{{\bm{w}}_{t-1}}}}\right](s_{t})}\right)$ $=2\iota_{K}{\bm{\phi}}(s_{t})\left({{\bm{p}}_{{\bm{w}}_{t-1}}(s_{t})-{\bm{p}}_{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}_{{\bm{w}}_{t-1}}}}(s_{t})}\right)^{\top}{\bm{C}}^{\top}{\bm{C}}$
$=2\iota_{K}{\bm{\phi}}(s_{t})\left[{{\bm{\phi}}(s_{t})^{\top}{\bm{W}}_{t-1}+\left({\frac{1}{K+1}\bm{1}_{K}-{\bm{p}}_{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}_{{\bm{w}}_{t-1}}}}(s_{t})}\right)^{\top}}\right]{\bm{C}}^{\top}{\bm{C}}.$